OpenStudy (anonymous):

Please reply if you can help me with slope and similar triangles? i will give a gold medal

OpenStudy (raden):

so, what's ur question then

OpenStudy (anonymous):

question 1

OpenStudy (raden):

obviously, it is 2/3 actually, it's called the slope

OpenStudy (anonymous):

can you help with more?

OpenStudy (raden):

sure

OpenStudy (anonymous):

number 2

OpenStudy (raden):

the slope of red triangle is 2/3, while for blue triangle is 4/6 = 2 * 2/3 it means twice of the slope of red triangle

OpenStudy (raden):

but, looks D is right too

OpenStudy (anonymous):

so what should i go with

OpenStudy (raden):

damned, i make big mistake! 4/6 = 2/3 so, they are be same, sorry

OpenStudy (anonymous):

its ok lol i have two more questions wanna help

OpenStudy (raden):

sure

OpenStudy (raden):

post it

OpenStudy (anonymous):

4

OpenStudy (anonymous):

5 the last one

OpenStudy (raden):

the formula to get the slope of line with passes through points (x1,x2) and (x2,y2) is m = (y2-y1)/(x2-x1) it is C

OpenStudy (anonymous):

ok and wb 5

OpenStudy (raden):

it is D the product of their slopes always be -1

OpenStudy (anonymous):

do you wanna help me with a quiz